OpenStudy (anonymous):

At a local plant sale, Janice was selling Golden Barrel Cactus plants and baby Saguaros. During the sale, she sold 5 times as many Saguaros ($6 each) as Golden Barrels ($10 each). She made a total of $720 during the plant sale. How many Saguaros did Janice sell? A. 18 B. 40 C. 72 D. 90

OpenStudy (aaronq):

lol just because its about plants it does not mean that it's biology related

OpenStudy (aaronq):

il help ya nonetheless

OpenStudy (aaronq):

basically use the info given to set up 2 equations: cactus =x saguaros =y (1) y=5x (2)6y+10x=720 substitute (1) into (2) 6(5x)+10x=720 solve for x then use (1) (knowing what x is) to solve for y.